a number of capacitors, each of capactiance 1uf and each one of which gets punctured if a potential difference just exceeding 500 volt is applied.are provided .Then an arrangement suitable for giving a capacitor of capacitance 3uf across which 2000 volt may be applied requirs at least

Grade:9

1 Answers

Khimraj
5 years ago
first for 2000V we have to require 2000/500 = 4 resistor in series.
equivalent of this = ¼ uF
for 3uF we will require 12 such series combination
Then no of capacitor = 12*4 = 48
Hope it clears.